Determine the horizontal and vertical components of a bus if it is moving at 80km/h going 22° south of west


1
Expert's answer
2021-10-01T12:17:33-0400


Let "\theta =22\degree" and "\mathbf{v}" be the vector of velocity with magnitude "v = 80km\/h".

Then its horizontal component is:


"v_x = -v\cos\theta = -80\cdot \cos22\degree \approx -74.2m/s"

Vertical component:


"v_x = -v\sin\theta = -80\cdot \sin22\degree \approx -30.0m/s"

Answer. Horiaontal: -74.2 m/s, vertical: -30.0 m/s.